# Threshold Graphs are Globally Synchronizing

### v2

@@ 2. Preliminaries of Threshold Graphs
Threshold graphs are a well-studied class of graphs that play a central role in graph theory and have found applications across various disciplines such as scheduling theory, psychology, computer science [threshold_graphs_book]. Threshold graphs have been defined by Chvs and Hammer as follows [ChvatalHammer1973, ChvatalHammer1977].
@@
The resulting graph on vertices $\{1,\dots,n\}$ is uniquely encoded by the binary sequence $(b_{2},\dots,b_{n})$, where $b_{i}=0$ if vertex $i$ is added as isolated and $b_{i}=1$ if it is added as dominating.
Although in this work we rely only on the construction process of threshold graphs to analyze their global synchronization, it is of independant interest to recall that they admit several equivalent characterizations [threshold_graphs_book].
@@
Remark 2.3 (Edge densities of threshold graphs).
The edge density of threshold graphs ranges from $O(1/n)$ to $1$. Indeed, if $(b_{2},\dots,b_{n})$ is the generating sequence, then
$|E|=\sum_{i:b_{i}=1}(i-1),\qquad\mathrm{dens}(G)=\frac{|E|}{\binom{n}{2}}$.
By choosing the positions at which $b_{i}=1$, one obtains connected threshold graphs whose densities range from $2/n$ (stars) to $1$ (the complete graph), spanning many intermediate values. Moreover, threshold graphs are exactly the graphs whose degree sequences saturate the Erdős–Gallai inequalities in the characteristic Ferrers pattern corresponding to nested neighborhoods [threshold_graphs_book].
@@
Interestingly, threshold graphs appeared in the synchronization literature long before their role in the Kuramoto model was considered. To the best of our knowledge, the earliest instance dates back to 1977, with the seminal work of Ausiello, Messina, and Protasi [Ausiello1977], who provided a graph-theoretic characterization of the PVc class of synchronizing primitives and introduced the notion of PVc-definable graphs. This line of work emerged almost simultaneously with the independant introduction of threshold graphs by Chvátal and Hammer in their study of aggregation of inequalities in integer programming [ChvatalHammer1977].
@@ 3.1. Equilibrium and Local Stability
By the definition, equilibrium points of a dynamical system are the points where the system remains at rest, meaning that the first-order derivative of the state vanishes [khalil2002nonlinear]. Setting the right-hand side of (2) to zero provides the characterization of an equilibrium point in the homogeneous Kuramoto model on the graph $G$.
@@
An equilibrium point is considered stable if small perturbations around it remain small over time [khalil2002nonlinear]. The linearization of a nonlinear system provides a necessary condition for local stability: an equilibrium point is stable if the Jacobian matrix is negative semi-definite.
@@ 3.3. Correspondence between Stability and Optimality
One can further relate the stability of an equilibrium to the second-order necessary condition of the energy landscape. It is precisely because of the real analycity of the energy function (4), the local stability of an equilibrium of (2) is equivalent to the local minimality of the energy function (4). The relation mentioned here is summarized in Figure 4. Such an equivalence was established in the celebrated work [Absil2006StableEquilibrium].
SOSP of 4 Local minima of 4 = stable equilibria of 2
@@ 4.2. Alignment and Equilibrium
The proof can be found in [monzon2005global], we present here for completeness.
@@ 4.3. Angle and Stability
where the second equality follows from the equilibrium condition (4.1), after rewriting the unit vectors in complex form. If some $\mu_{i}<0$, then $\mathrm{Hess}\,E(\theta)_{ii}<0$, contradicting the positive semidefiniteness of $\mathrm{Hess}\,E(\theta)$ at any second-order stationary point [Bhatia2007] [HornJohnson1985]. ∎
@@ 6.1. A Prototype Example
7. Discussion and Open Problems
@@
In this work, we present a class of globally synchronizing graphs — threshold graphs — whose density vary from $2/n$ to $1$, revealing a new mechanism of global synchronization that has not been articulated before. Our proof crucially exploits the constructive nature of threshold graphs: they form the smallest graph class closed under the addition of isolated nodes and universal nodes. This closure property is essential to our inductive argument, as it allows us to show that global synchronization is preserved when starting from a threshold graph and adding either an isolated node or a universal node.
@@
Moreover, we expect that our results may admit further generalizations. Possible directions include extending the analysis to oscillators evolving on higher-dimensional spheres, incorporating nonlinear interaction terms as in [Geshkovski2025], and considering broader classes of graphs. For instance, one may study systems in which oscillators are coupled through both positive and negative edges.
@@ Acknowledgements
We would like to thank Afonso S. Bandeira for insightful discussions.

### v4

@@ 2. Preliminaries of Threshold Graphs
Threshold graphs are a well-studied class of graphs that play a central role in graph theory and have found applications across a wide range of disciplines. They were discovered independently by researchers working in different fields and reported in separate venues. The term threshold graph was introduced by Chvátal and Hammer [CH73, CH77] in their study of set-packing problems. Around the same time, the same class of graphs was independently identified by Henderson and Zalcstein in the context of parallel processing, where they were referred to as PV-chunk definable graphs. Since then, threshold graphs have appeared in numerous settings, including resource allocation, cyclic scheduling, manpower planning, complexity theory, and the study of graphical degree sequences [MP95].
2.1. Definition and Characterizations
@@
The resulting graph on vertices $\{1,\dots,n\}$ is uniquely encoded by the binary sequence $(b_{2},\dots,b_{n})$, where $b_{i}=0$ if vertex $i$ is added as isolated and $b_{i}=1$ if it is added as dominating node.
Although in this work we rely only on the construction process of threshold graphs to analyze their global synchronization, it is of independant interest to recall that they admit several equivalent characterizations [MP95].
@@
Interestingly, threshold graphs appeared in the synchronization literature long before their role in the Kuramoto model was considered. To the best of our knowledge, the earliest instance dates back to 1977, with the seminal work of Ausiello, Messina, and Protasi [AMP77], who provided a graph-theoretic characterization of the PVc class of synchronizing primitives and introduced the notion of PVc-definable graphs. This line of work emerged almost simultaneously with the independant introduction of threshold graphs by Chvátal and Hammer in their study of aggregation of inequalities in integer programming [CH77].
@@ 3.1. Equilibrium and Local Stability
By the definition, equilibrium points of a dynamical system are the points where the system remains at rest, meaning that the first-order derivative of the state vanishes [Kha02]. Setting the right-hand side of (2) to zero provides the characterization of an equilibrium point in the homogeneous Kuramoto model on the graph $G$.
@@
An equilibrium point is considered stable if small perturbations around it remain small over time [Kha02]. The linearization of a nonlinear system provides a necessary condition for local stability: an equilibrium point is stable if the Jacobian matrix is negative semi-definite.
@@ 3.3. Correspondence between Stability and Optimality
One can further relate the stability of an equilibrium to the second-order necessary condition of the energy landscape. It is precisely because of the real analycity of the energy function (4), the local stability of an equilibrium of (2) is equivalent to the local minimality of the energy function (4). The relation mentioned here is summarized in Figure 4. Such an equivalence was established in the celebrated work [AK06].
SOSP of (4) Local minima of (4) = stable equilibria of (2)
@@ 4.2. Alignment and Equilibrium
where $v_{i}:=(\cos\theta_{i},\sin\theta_{i})$.
The proof can be found in [MP05], we present here for completeness.
@@ 4.3. Angle and Stability
where the second equality follows from the equilibrium condition (4.1), after rewriting the unit vectors in complex form. If some $\mu_{i}<0$, then $\mathrm{Hess}\,E(\theta)_{ii}<0$, contradicting the positive semidefiniteness of $\mathrm{Hess}\,E(\theta)$ at any second-order stationary point [Bha07] [HJ85]. ∎
@@ 6.1. A Prototype Example
7. Discussion: Beyond Density-Based Guarantees for Global Synchronization
As noted in the introduction, previous guarantees for global synchronization have largely focused on minimum-degree conditions: that is, requiring every node to be connected to more than a fixed fraction ($\geq 75\%$) of the remaining vertices. In this regime, graphs are necessarily highly dense, and their degree sequences are relatively homogeneous. This naturally raises the question:
Does a graph that exhibits global synchronization necessarily have high edge density or a homogeneous degree sequence (i.e., with relatively small variation in node degrees)?
In this section, we provide a negative answer to this question [^5] Of course, trees provide trivial examples. by exploiting two structural properties of threshold graphs, which were shown in the previous sections to be globally synchronizing. More precisely, we show that threshold graphs realize every admissible edge density among connected graphs. Moreover, their degree sequences are highly heterogeneous: among all connected graphs with a fixed edge density, threshold graphs attain the maximally left-skewed degree sequence when ordered in nonincreasing order.
7.1. Threshold Graphs Realize All Densities
Note that the edge density of a graph $G$ on $n$ vertices takes values in
$\mathrm{density}(G)\in\left\{\,\frac{m}{\binom{n}{2}}\;\middle|\;m\in\{0,\ldots,\binom{n}{2}\}\,\right\}$. (20)
and for connected graphs
$\mathrm{density}(G)\in\left\{\,\frac{m}{\binom{n}{2}}\;\middle|\;m\in\{n-1,\ldots,\binom{n}{2}\}\,\right\}$. (21)
If the graph is not connected, to realize any density in $\left\{\,\frac{m}{\binom{n}{2}}\;\middle|\;m\in\{0,\ldots,n-2\}\,\right\}$, construct a threshold graph on $n$ vertices whose n1n-1-length sequence contains $1$ at the $m$-th position.
001
Figure 15. Illustration of the threshold graph construction. A single 11 at the third position of the generating sequence adds a dominating vertex, which contributes three incident edges shown as curved arcs.
The following lemma shows that every admissible edge density in the range (21) is realized by a connected threshold graph on $n$ vertices.
Lemma 7.1.
Fix $n\geq 2$. For any integer $m\in[\,n-1,\binom{n}{2}\,$], there exists a connected threshold graph on $n$ vertices with exactly $m$ edges.
Proof.
Since the graph is connected, the last vertex added in the construction of a threshold graph must be a dominating vertex, contributing n1n-1 edges. To obtain a graph with exactly $m$ edges, the vertices added before the last one must therefore contribute a total of m(n1)m-(n-1) edges. Since $m\in\{n-1,\ldots,\binom{n}{2}\}$, it follows that
$m-(n-1)\in\left\{0,1,\ldots,\binom{n}{2}-(n-1)\right\}=\left\{0,1,\ldots,\frac{(n-2)(n-1)}{2}\right\}.$
Edges are created only when dominating vertices are added during the construction. We fix the following convention: starting from an initial vertex, we index the construction sequence by counting only the dominating vertices added thereafter. Under this convention, a dominating vertex added at position $p$ contributes exactly $p$ new edges, as it connects to all previously existing vertices. Hence the possible edge contributions of dominating vertices, except for the last one added in the construction sequence, take values in $\{1,2,\ldots,n-2\}$.
Therefore, the problem reduces to determining at which positions the dominating vertices except for the last one are added so that they collectively contribute m(n1)m-(n-1) edges. More precisely, it suffices to prove the following combinatorial statement:
For any integer
$k\in\left\{0,1,\ldots,\frac{(n-2)(n-1)}{2}\right\},$
there exists a subset
$A\subseteq\{1,2,\ldots,n-2\}\quad\textit{such that}\quad\sum_{a\in A}a=k.$
We prove this claim by induction on $n$.
Base case. For $n=3$, we have $\{1,2,\ldots,n-2\}=\{1\}$ and $k\in\{0,1\}$. If $k=0$, take $A=\emptyset$; if $k=1$, take $A=\{1\}$.
Induction hypothesis. Assume that for any
$k\in\left\{0,1,2,\ldots,\frac{(n-3)(n-2)}{2}\right\},$
there exists a subset
$A^{\prime}_{k}\subseteq\{1,2,\ldots,n-3\}\quad\text{such that}\quad\sum_{a\in A^{\prime}_{k}}a=k.$
Induction step. We now consider any
$k\in\left\{0,1,2,\ldots,\frac{(n-2)(n-1)}{2}\right\}.$
Case 1. If
$k\leq\frac{(n-3)(n-2)}{2},$
then by the induction hypothesis there exists $A^{\prime}_{k}\subseteq\{1,2,\ldots,n-3\}$ such that $\sum_{a\in A^{\prime}_{k}}a=k$. Setting $A:=A^{\prime}_{k}$ yields the desired subset.
Case 2. If
$k>\frac{(n-3)(n-2)}{2},$
then using only values in $\{1,2,\ldots,n-3\}$ cannot achieve the sum $k$, since the maximal possible sum is
$\sum_{i=1}^{n-3}i=\frac{(n-2)(n-3)}{2},$
which is strictly smaller than $k$. Hence the element n2n-2 must be included in $A$.
It then remains to find a subset
$A^{\prime\prime}\subseteq\{1,2,\ldots,n-3\}\quad\text{such that}\quad\sum_{a\in A^{\prime\prime}}a=k-(n-2).$
Since
$k-(n-2)\in\left\{\frac{(n-2)(n-5)}{2}+1,\ldots,\frac{(n-2)(n-3)}{2}\right\}\subseteq\left\{0,1,\ldots,\frac{(n-3)(n-2)}{2}\right\},$
the induction hypothesis guarantees the existence of such a subset; in particular, we may take
$A^{\prime\prime}=A^{\prime}_{\,k-(n-2)}.$
Setting $A:=A^{\prime\prime}\cup\{n-2\}$ then yields $\sum_{a\in A}a=k$.
This completes the induction and hence the proof. ∎
7.2. Degree Sequence of Threshold Graphs
One way to compare the “shape” of the degree sequences of two graphs with the same edge density is through majorization, which provides a partial order quantifying how spread out or concentrated the degrees are. More specifically, let $G$ and $G^{\prime}$ be two graphs with the same edge density, and let
$\mathbf{d}(G)=(d_{1},\dots,d_{n}),\qquad\mathbf{d}(G^{\prime})=(d^{\prime}_{1},\dots,d^{\prime}_{n})$
denote their degree sequences arranged in nonincreasing order. We say that $\mathbf{d}(G)$ majorizes $\mathbf{d}(G^{\prime})$, and write $\mathbf{d}(G)\succ\mathbf{d}(G^{\prime})$, if
$\sum_{i=1}^{k}d_{i}\geq\sum_{i=1}^{k}d^{\prime}_{i}\quad\text{for all }k=1,\dots,n-1,$
and
$\sum_{i=1}^{n}d_{i}=\sum_{i=1}^{n}d^{\prime}_{i}.$
The following is a classical result on degree sequence (called a threshold sequence) of threshold graphs.
Theorem 7.2 ([PS89]).
A degree sequence $\mathbf{d}$ is a threshold sequence if and only if it is not strictly majorized by any other graphical degree sequence.
As shown in the previous section, for any admissible edge density there exists at least one (not necessarily unique) threshold graph realizing that density. Moreover, Theorem 7.2 tells that, among all graphs realizing a given edge density, threshold graphs are exactly those whose degree sequences are maximal with respect to majorization. In this sense, threshold graphs are always the most left-skewed among graphs with a fixed edge density, since their degree sequences are maximal under majorization, meaning that the degree mass is as concentrated as possible on high-degree vertices.
8. Summary and Open Problems
@@
In this work, we present a class of globally synchronizing graphs — threshold graphs — whose density vary from $2/n$ to $1$. Our proof crucially exploits the constructive nature of threshold graphs: they form the smallest graph class closed under the addition of isolated nodes and universal nodes. This closure property is essential to our inductive argument, as it allows us to show that global synchronization is preserved when starting from a threshold graph and adding either an isolated node or a universal node.
@@
Moreover, we expect that our results may admit further generalizations. Possible directions include extending the analysis to oscillators evolving on higher-dimensional spheres, incorporating nonlinear interaction terms as in [GLPR25], and considering broader classes of graphs. For instance, one may study systems in which oscillators are coupled through both positive and negative edges.
@@ Acknowledgements
We would like to thank Afonso S. Bandeira for insightful discussions and Kevin Lucca for pointing out relevant references on the known bounds of the density threshold.